Harry drove for 2 hours on the​ freeway, then decreased his speed by 20 miles per hour and drove for 5 more hours on a country road. If his total trip was 271 ​miles, then what was his speed on the​ freeway?
Answer by ikleyn(52865)

.

Let x be his speed on the​ freeway, in miles per hour.


Then his speed on the country road was (x-20) mph.


The full distance equation


    2x + 5*(x-20) = 271.


    2x + 5x = 271 + 100

    7x      = 371

     x      = 371/7 = 53.


ANSWER.  His speed on the​ freeway was 53 mph.

Solved.

----------------

See introductory lessons on Travel and Distance problems
    - Travel and Distance problems
    - Travel and Distance problems for two bodies moving in opposite directions
    - Travel and Distance problems for two bodies moving in the same direction (catching up)
in this site.

They are written specially for you.

You will find the solutions of many similar problems there.

Read them and learn once and for all from these lessons on how to solve simple Travel and Distance problems.

Become an expert in this area.